# Chamani Shiranthika, Zahra Hafezi Kafshgari, Hadi Hadizadeh, Parvaneh Saeedi School of Engineering Science

### v1

@@ 1 Introduction
The structure of this paper is as follows: Section 2 provides an overview of SplitFed fundamentals and the current state-of-the-art in SplitFed for medical image segmentation, including its integration with emerging transformer architectures. It also discusses publicly accessible federated network repositories and offers insights into semantic segmentation and the architecture of our implemented split models, followed by a discussion on our split point selection decisions. Section 2.5 details the experiments, results, and evaluations. Section 3 highlights the potential future directions. This paper concludes by presenting the conclusions in Section 4.
@@ 2 Related works
This section reviews related work on split federated learning, publicly available federated repositories, and semantic segmentation, including notable semantic segmentation networks. Also, it offers insights into our decisions on split point selection for designing our split networks.
@@ 2.1 Split Federated Learning (SplitFed)
SplitFed is the combination of FL and SL[5]. In FL, a central server collaboratively builds a global model using data from multiple devices or clients without accessing sensitive local information [21, 22]. During collaborative training, clients are tasked with training the entire model on their respective local platforms. Consequently, some clients encounter challenges, particularly in comparison to local training, stemming from limited computational capabilities or resource-constrained issues. Split learning (SL) overcomes this issue by splitting and residing the model parts at different locations, devices, or clients [23]. While federated structures protect client data privacy by securely storing raw information on local devices, splitting the model further alleviates potential computational burdens on clients. Therefore, SplitFed learning is particularly advantageous in scenarios where data cannot be centralized for any reason. It strikes a balance between collaborative model training with limited computational resources, and preserving the privacy of individual user data.
@@ 2.2 Publicly available federated repositories
Despite the availability of these resources, our survey reveals a gap in the current literature: the absence of a reusable split federated network repository that enables comprehensive comparisons and benchmarks. This gap highlights an opportunity for future research and development to create a robust and adaptable split network repository that supports collaborative networks. Our focus is on addressing this need by facilitating and streamlining the training processes for split federated networks in the healthcare domain.
@@ 3.0.1 Transformers in SplitFed
3.0.2 Design of cross-domain architectures
@@
3.0.3 Personalized network architectures
@@
3.0.4 Privacy-preserving architectures
@@
3.0.5 Communication-efficient architectures
@@
3.0.6 Design of SplitFed-specific network architectures

### v2

@@ 1 Introduction
The structure of this paper is as follows: Section 2 provides an overview of SplitFed fundamentals and the current state-of-the-art in SplitFed for medical image segmentation, including its integration with emerging transformer architectures. It also discusses publicly accessible federated network repositories and offers insights into semantic segmentation and the architecture of our implemented split models, followed by a discussion on our split point selection decisions. Section 3 details the experiments, results, and evaluations. Section 4 highlights the potential future directions. This paper concludes by presenting the conclusions in Section 5.
@@ 2 Related works
This section reviews related work on SplitFed learning, publicly available federated repositories, and semantic segmentation, including notable semantic segmentation networks. Also, it offers insights into our decisions on split point selection for designing our split networks.
@@ 2.1 Split Federated Learning (SplitFed)
SplitFed is the combination of FL and SL[5]. In FL, a central server collaboratively builds a global model using data from multiple devices or clients without accessing sensitive local information [21, 22]. During collaborative training, clients are tasked with training the entire model on their respective local platforms. Consequently, some clients encounter challenges, particularly in comparison to local training, stemming from limited computational capabilities or resource-constrained issues. SL overcomes this issue by splitting and residing the model parts at different locations, devices, or clients [23]. While federated structures protect client data privacy by securely storing raw information on local devices, splitting the model further alleviates potential computational burdens on clients. Therefore, SplitFed learning is particularly advantageous in scenarios where data cannot be centralized for any reason. It strikes a balance between collaborative model training with limited computational resources, and preserving the privacy of individual user data.
@@ 2.2 Publicly Available Federated Repositories
Despite the availability of these resources, our survey reveals a gap in the current literature: the absence of a reusable SplitFed network repository that enables comprehensive comparisons and benchmarks. This gap highlights an opportunity for future research and development to create a robust and adaptable split network repository that supports collaborative networks. Our focus is on addressing this need by facilitating and streamlining the training processes for SplitFed networks in the healthcare domain.
@@ 2.3 Image Segmentation Models
  34. 34.
    SAM[82]: Segment Anything
  35. 35.
    VM-UNet[83]: Vision Mamba UNet
  36. 36.
    HC-Mamba[84]: Hybrid-convolution version of Vision Mamba
  37. 37.
    EoMT[85]: Encoder-only Mask Transformer
  38. 38.
    Med-SA[86]: Medical SAM Adapter
@@ 4.1 Transformers in SplitFed
4.2 Design of Cross-Domain Architectures
@@
4.3 Personalized Network Architectures
@@
4.4 Privacy-Preserving Architectures
@@
4.5 Communication-Efficient Architectures
@@
4.6 Design of SplitFed-Specific Network Architectures